OpenStudy (help_people):

Given the following functions f(x) and g(x), solve (f ⋅ g)(2) and select the correct answer below: f(x) = 2x2 − 10 g(x) = x + 9 −22 22 −66 66

OpenStudy (help_people):

@Nurali

OpenStudy (nurali):

f(2)=8-10 f(2)=-2 g(2)=2+9 g(2)=11 (f.g)(2)=-2*11 (f.g)(2)=-22

OpenStudy (help_people):

thank you so much is it possible if you could help me w. one more > @Nurali

OpenStudy (nurali):

i will try

OpenStudy (help_people):

well actually 2

OpenStudy (help_people):

possibly would you like me to repost on to another tab or ?

OpenStudy (help_people):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 2x + 21 g(x) = x − 24 −27 −21 6 48

OpenStudy (nurali):

f(x)+g(x) 2x+21+x-24 3x-3 3(3)-3 9-3 6

OpenStudy (help_people):

tha k you one more pretty please ?

OpenStudy (help_people):

@Nurali

OpenStudy (nurali):

Try

OpenStudy (help_people):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below. m[n(x)] = 4x − 51 m[n(x)] = 4x − 29 m[n(x)] = 4x2− 51 m[n(x)] = 4x2 − 29

OpenStudy (help_people):

@Nurali ^

OpenStudy (nurali):

4x-40-11 4x-29

OpenStudy (help_people):

thank you :)

OpenStudy (help_people):

i got that one wrong :(